# Analysis of optical pattern formation on glass

### v1

@@ 3.4 Engineered patterns
We were able to engineer the two desired patterns, a square and a circle, using the specular holography method (see Fig. 11); the video showing the complete process for hologram creation is available as supplementary material.
@@
Figure 11: Two patterns created by scratching the CD cases: (a) a square hologram and (b) a circle hologram.
@@ 3.5 Further studies
One way to improve this study is by verifying the equivalency between reflected and transmitted patterns. We used this fact to analyze only the reflected one in the simulation, but to certify it in the experiments one would need to capture the halo at the same position and angle on both sides of the surface.
Another difficulty is faced while observing the effects of distance on the halo density. The point of light incidence on the glass needs to be accurate every time we decrease the distance, ensuring that it is the only varying factor. While we tried to maintain this caution in the experiments, one would need a more precise experimental setup to ensure the results and compare the densities.
@@ 4 Conclusions
We investigated the formation of optical patterns on glass. We observed experimentally that randomly scratched glass plates produce the glass halo phenomenon, highlighting the structural inhomogeneities. The second pattern we studied was the formation of diverging lines, which comes from an optical distortion of the imaging process; indeed we observed that they rotate as we rotate the camera. We assumed that the phenomenon of light passing through a glass that has been mechanically scratched can be acceptably described using ray optics; indeed a simulation based on this assumption well reproduced our experiments. According to our results, the halo formation on everyday reflective objects, such as glass, is governed by specular reflection on various points along the scratches. We used our simulation to study the effect of two factors on the halo density, namely, (1) the number of scratches and (2) the distance between the surface and the camera-light source. As we progressively increase the number of scratches on the surface, the halo density also increases; similarly, we see a fuller circle of halos when the plate is relatively far from the light-camera position.

### v2

@@ 3.4 Engineered patterns
We were able to engineer the two desired patterns, a square and a circle, using the specular holography method (see Fig. 12); the video showing the complete process for hologram creation and movement is available as supplementary material. The interesting feature of this method is that we can engineer any 3D shape or pattern by creating only circular scratches on the surface, provided that we follow the desired shape on a 2D blueprint.
@@
Figure 12: Two patterns created by scratching the CD cases: (a) a square hologram and (b) a circle hologram. The movement and 3D perspective is better seen in the available video.
@@ 3.5 Further studies
One way to improve this study is by verifying the equivalency between reflected and transmitted patterns. We used this fact to analyze only the reflected one in the simulation, but to certify it in the experiments one would need to capture the halo at the same position and angle on both sides of the surface, needing a very precise alignment mechanism.
It is extremely difficult to obtain quantitative analysis for the halo density. First, =we were not able to experimentally count the number of scratches to later compare the densities. Also, the point of light incidence on the glass needs to be accurate every time we decrease the distance, ensuring that it is the only varying factor. While we tried to maintain this caution in the experiments, one would need a more precise experimental setup to ensure the results and compare the halo density. Thus, we validated our simulation and use it to perform these analysis.
@@
3.6 Pedagogical application
Our goal so far was to present the physics inside the IPT competition. We have shown the steps taken as we tackled the problem and our solution.
Now we discuss its didactic use. First, the experimentation with scratched glass plates is easy and can be conducted to teach the principles of geometric optics since the scratches can reflect light in all directions, causing the circular halo pattern. This particular pattern can be seen in various objects of everyday life, as seen in Fig. 1, and can be easily reproduced.
Also, the holography technique explored in this article shows how we can use simple tools to design optical patterns. In this context, the curved scratches act as spherical mirror, forming real or virtual images that we perceive as 3D images. Thus, these simple experiments communicate scientific knowledge related to reflection and image formation, being suitable for scientific fairs and didactic demonstrations in class.
Lastly, our simulation can act as a first contact with computational methods in physics for high school and undergraduate students. The simulation, written in Python, is accessible and exemplifies the use of arrays, functions and graphs. After understanding the principle of geometric optics, the student can use the simulation to analyze the role of each parameter in the optical outcome while learning a programming language.
@@ 4 Conclusions
We investigated the formation of optical patterns on glass. We observed experimentally that randomly scratched glass plates produce the glass halo phenomenon, highlighting the structural inhomogeneities. The second pattern we studied was the formation of diverging lines, which comes from a known optical distortion in the imaging process; indeed we observed that they rotate as we rotate the camera. We assumed that the phenomenon of light passing through a glass that has been mechanically scratched can be acceptably described using ray optics; indeed a simulation based on this assumption well reproduced our experiments. According to our results, the halo formation on everyday reflective objects, such as glass, is governed by specular reflection on various points along the scratches. We used our simulation to study the effect of two factors on the halo density, namely, (1) the number of scratches and (2) the distance between the surface and the camera-light source. As we progressively increase the number of scratches on the surface, the halo density also increases; similarly, we see a fuller circle of halos when the plate is relatively far from the light-camera position. Our methods can be used as a simple experimental demonstration of optical concepts in classes and science communication activities.
@@ Acknowledgements
Appendix
A.1 Simulation
In order to simulate the phenomenon, we follow [20]: suppose we have a reflective surface with a scratch whose curve can be parametrized by $\vec{r}(t)$ and whose unit tangent vector is $\hat{q}(t)$. If we have an arrangement such as in Fig. 3, then a point in the scratch will be seen if the following equation is true:
$\hat{E}(t)\cdot\hat{q}(t)=\hat{R}(t)\cdot\hat{q}(t),$ (8)
where $\vec{E}(t)=\vec{r}(t)-\vec{L}$, $\vec{R}(t)=\vec{C}-\vec{r}(t)$, and $\hat{E}(t)$ and $\hat{R}(t)$ represent their respective normalized vectors. One key assumption that has to be made is that the scratches have a cross sectional shape that is able to reflect at any angle, e.g. a shallow parabola or a shallow semicircle as seen in Fig. 13; if this was not assumed, then additional conditions would need to be established in order to guarantee that the camera is inside of the region where the light reflected off the scratch could reach it. Another assumption needed is that the light source is far away and point-like, to guarantee that the light rays are parallel and reflect off the scratch as seen in Fig. 13.
For a generic scratch curve, equation (8) is not able to be solved analytically, and therefore needs root finding algorithms that naturally introduce a tolerance for the equation. Numerically, what the computer does is finding values of $t$ that satisfy:
$\left|\hat{E}(t)\cdot\hat{q}(t)-\hat{R}(t)\cdot\hat{q}(t)\right|\leq TOL,$ (9)
or, calling $\theta_{1}$ the angle between $\hat{E}(t)$ and $\hat{q}(t)$ and, similarly, $\theta_{2}$ the angle between $\hat{R}(t)$ and $\hat{q}(t)$, the inequality in (9) can be written as
[…]
Since $|\hat{E}(t)|=1$, $|\hat{R}(t)|=1$ and $|\hat{q}(t)|=1$, the camera or our eyes will capture the reflected beam if the following relation is satisfied, within a narrow tolerance,
$cos(\theta_{1})-cos(\theta_{2})\leq TOL.$ (11)
Figure 13: Schematic representation of the assumed scratch cross-sectional shape. The cross section of the scratch is assumed to be one such that rays from the light source (coming from above in the image) can reflect at any angle. The shape used to make the above image was a shallow parabolic mirror obtained using [32].
To illustrate the usage of equation (5), we consider a the simplest example that can be solved analytically: a line. A line $\vec{r}(t)$ can be written as
$\vec{r}(t)=\vec{P}_{o}+\hat{V}t,$ (12)
where $\vec{P}_{o}$ is the “origin” of the line, i.e. the point where $t=0$, and $\hat{V}$ is the direction in which the line points with increasing $t$, which is also the tangent vector to the line. We will, without loss of generality, assume that $\hat{V}$ is normalized as notation implies.
Let us now consider the light source and the camera to be at symmetrical positions with respect to the $yz$ plane, that is, if $\vec{C}=(d,0,h)$, then $\vec{L}=(-d,0,h)$. With these vectors defined, we can use equation (5) to get
$\vec{r}(t)\cdot\hat{V}=\left(\frac{\vec{C}+\vec{L}}{2}\right)\cdot\hat{V}$. (13)
As $\vec{C}$ and $\vec{L}$ are symmetrical, then $(\vec{C}+\vec{L})/2=(0,0,h)$. Furthermore, since $\hat{V}$ is a vector parallel to the $xy$ plane, its $z$ component is zero, which makes the right hand side of the equation equal to zero. We now expand the left hand side using the definition of $\vec{r}(t)$:
$\left(\vec{P}_{o}+\hat{V}t\right)\cdot\hat{V}=0.$ (14)
Therefore, the value of $t$ that satisfies this equation is
$t=-\vec{P}_{o}\cdot\hat{V}$, (15)
and the point along the line that would be seen because of the reflection would be
$\vec{r}(t)=\vec{P}_{o}-\left(\vec{P}_{o}\cdot\hat{V}\right)\hat{V}$. (16)
For example, if the line was such that it passes through the point $\vec{P}_{o}=(1\mbox{\,m},0,0)$ and has a slope of $45^{\circ}$ ($\hat{V}=1/\sqrt{2}(1\mbox{ m},1\mbox{ m},0)$) with respect to the $x$ axis, then the point on this line that would be seen is $1/2(1\mbox{ m},-1\mbox{ m})$.
One benefit of the simulation is that it “automatically” accounts for the practical aspects of our retinas or the cameras’ detectors. In practice, we observe not singular points, but certain segments that are illuminated from each scratch; and that is because light rays that are close enough to the point calculated to be the position of the glint also manage to reach the eye or camera because of the non-zero size of their apertures. Therefore, even for linear scratches it is better to simulate the phenomenon using equation (11), with a tolerance, rather than using the analytical solution obtained.
A.2 Scratch reconstruction
As explained in Sec. 3.2, the reconstruction using the Defocusing Microscopy technique reveals how much the scratches vary in size; the standard deviation in Table 3 confirms that they are on the same scale (micrometers) but can differ significantly, since the sandpaper is not completely uniform and the grains can produce scratches of different sizes. We have shown the width distribution for all scratches used in our analysis (Fig. 7). We highlight that they are larger than the wavelength of light and are concentrated around $18\,\mu$ m. Lastly, in Fig. 14 we present the distribution of the scratch width and depth for plates 4, 5 and 6.
Figure 14: Distribution of the width and depth size for the randomly scratched plates.